# RotBench: Evaluating Multimodal Large Language Models on Identifying Image Rotation

### v2

@@ 6.1 Model Bias Towards 0° and 90°
Figure 4: Confusion matrix of true vs. predicted rotations for GPT-4o using CoT prompting, summed across three runs on RotBench-large. Rows represent ground-truth labels, columns represent predicted labels.
To further elucidate the specific types of rotational errors made by models, we analyze the confusion matrix for GPT-4o. Figure˜4 shows the confusion matrix obtained from summing predictions across three runs. We see GPT-4o predominantly struggles with differentiating between 90° and 270° rotations. Specifically, the model misclassifies 459 instances of 90° images as 270° and 424 instances of 270° images as 90°. Yet, 0° and 180° show significantly fewer misclassifications. This analysis underscores a critical shortcoming in model performance, suggesting either that the vision encoder is not providing sufficient signals to distinguish between clockwise and counter-clockwise rotations, or that the MLLM is not adequately incorporating the visual information into its reasoning.
@@ 6.2 Distinguishing Clockwise from Counter-clockwise Rotations
Table 3: Accuracy of different models in identifying 90° clockwise (CW) versus counter-clockwise rotation (CCW). Column indicates ground truth rotation (GT). A small number of responses are incorrectly identified as being right-side up. These responses are excluded in the table’s statistics.
@@
Following the analysis in Section˜6.1, we specifically examine whether MLLMs can distinguish clockwise (CW) and counter-clockwise rotations (CCW). For this, we simplified the original four-way classification task into a binary classification task. Using the 90° and 270° images from RotBench-large, we ask MLLMs to determine whether an image has been rotated 90° CCW or CW. [^6] Note that 90° CW is equivalent to a 270° CCW. By asking models to explicitly differentiate the two directions of rotation, we hope to provide a clearer signal for directional understanding. Figure˜5 shows an example question where the GPT-4o provides an incorrect answer.
Results.Table˜3 tests CW vs. CCW classification on GPT-4o and Qwen-2.5-VL-7B-Instruct (we additionally provide similar smaller-scale tests on RotBench-small for GPT-5 in Appendix˜E). Table˜3 indicates that GPT-4o has a significant bias toward identifying rotations as 90° counter-clockwise. GPT-4o correctly identified only 52 out of 300 clockwise rotations, whereas Qwen-2.5-VL-7B-Instruct correctly identified only 23 out of 300 clockwise rotations. The models consistently default to labeling ambiguous or uncertain rotations as counter-clockwise, hinting towards a potential underlying perceptual bias (Appendix˜F). While we find that the most recent model tested, GPT-5, has less bias and is better able to identify clockwise rotations, this improvement is not reflected in Table˜2, where the model performs on par with others; in other words, poor performance on RotBench cannot be solely explained by a model’s ability to distinguish clockwise from counterclockwise rotation. See Appendix˜E for further discussion.
These findings strongly indicate that the MLLMs tested have limitations in reliably distinguishing between CW and CCW rotational directions, offering a potential explanation for why distinguishing 90° and 270° CCW rotations presents such a challenging task. Further corroborating this idea, we show that evaluating Qwen-2.5-VL-7B-Instruct, GPT-4o, o3 on RotBench-small using clockwise angles does not result in significant performance differences (Appendix˜D).
@@ 6.3 Can Fine-tuning Solve RotBench?
Setup.  To assess whether specialized training can mitigate these performance issues, we conduct fine-tuning experiments using Qwen-2.5-VL-7B-Instruct. During data filtering, we find some images in Spatial-MM closely resemble each other (Fig.˜11). If we fine-tune on images in Spatial-MM not selected for RotBench, the existence of similar images in the training and testing sets may lead to inflated performance. To prevent such instances of overfitting, we train on 1000 images from MS COCO Lin et al. (2015) and evaluate performance on RotBench-large. Appendix˜B provides further training details.
Results.Fig.˜6 reveals a high and consistent accuracy for 0° throughout training, indicating robust recognition of upright images. Performance on 180° gradually improves, stabilizing around 0.8 after approximately 7000 images. However, accuracies for 90° and 270° exhibit substantial oscillations, suggesting that the model did not achieve stable improvement. The model appears to be caught in a cycle, alternating between accuracy gains and losses for these two rotations. This phenomenon is also reflected in our main results (Section˜5), where using CoT prompting improved accuracy on 270° images at the expense of 90°. The unstable performance may result from potential representational constraints in current visual encoders that limit visual understanding capabilities, particularly regarding subtle rotational distinctions.
@@ A.1 Further details on Stage 1 and 2 Filtering
Stage 1.  We randomly sample 300 images from Spatial-MM Shiri et al. (2024). Spatial-MM is divided into two splits, one-subject and two-subject, respectively comprised of images that contain one and two primary subjects. As models may use image dimensions to infer orientation, we crop each image into a square before further processing. We sample 150 images from each split and show the resulting 300-image dataset to the Stage 1 annotator. For each image, the annotator can decide to either accept, discard, or flag. The annotator accepts an image if it (1) contains easily identifiable rotational visual cues (e.g., a person standing), and (2) the image has meaningful differences when rotated. Images that do not satisfy these criteria are discarded. Occasionally, some images have subtle visual cues, or require more detailed semantic understanding to correctly perceive orientation. These images tend to have primary subjects that do not vary significantly with rotation, or require integrating background signals to identify rotation. The sample image shown in Fig.˜7 is an example of a flagged image. Such images are flagged by the Stage 1 annotator to proceed to Stage 2. We provide further examples and details of accepted, flagged, and discarded images in Fig.˜9. Note that the Stage 1 annotator is also tasked with ensuring no leaked personal information or offensive content is incorporated into RotBench.
@@
Of the first 300 sampled images, only 27 were flagged. Among the flagged images, only two are eventually rejected. We sample two additional images to replace the discarded images. These images form the 300-image RotBench-large benchmark.
@@ A.4 Annotator Interface
Figure˜8 shows the interface provided to Stage 2 annotators. The mapping between letter choice and rotation degree is shuffled from question to question.
@@ A.5 Sample Images
Figure˜3 displays the various forms of additional information provided to the model. Figure˜9 provides examples of an image that is accepted, flagged, and discarded during Stage 1 of data filtering. The accepted image has a clear subject and is easily distinguishable when rotated 0°, 90°, 180°, and 270°. The flagged image is more difficult to identify, however, the slight tilt – as opposed to a directly top-down perspective – still enables accurate judgment of rotation. However, the discarded image has no meaningful signals to distinguish between the four orientations.
@@ A.6 Further details on flagged images
Flagged images typically involve cases where there is a lack of primary subjects or the primary subject does not offer any distinguishable signals under rotation. These images tend to require deeper comprehension in order to make an accurate classification. For instance, Fig.˜7 displays a blank projector canvas above some colorful couches on a patio. The orientation of a projector screen cannot be reliably identified by where its supporting structures are located – the screen can be hung (cable hangs down), pitched up (supporting stand on the ground), or attached to a wall (support frame located sideways of the screen). Therefore, correctly identifying the orientation of the image requires understanding the correct orientation of the couches.
Appendix B Model Inference and Training Details
@@
Appendix C Additional Results
We provide experimental results from evaluating Llama-3.2-11B-Instruct on RotBench-large (Table˜4), as well as o3 and Gemini-2.5-Pro on RotBench-small (Table˜5). Llama-3.2-11B-Instruct seems to be the weakest model tested, showing inferior performance across all rotations. Without CoT, the model only achieves roughly 0.3 accuracy. The two frontier reasoning models (o3 and Gemini-2.5-Pro) show no significant performance improvement when prompted with auxiliary information. However, they respond positively to rotation grids. Gemini-2.5-Pro achieves the highest accuracy on 90° and 270° (both 0.7) using guided rotation grids, while o3 achieves its highest accuracy on 180° (0.83) using the rotation grid. These results indicate that strong reasoning models are able to utilize the grid to aid them in making a prediction, while weaker models struggle with it.
@@ Appendix C Additional Results
Appendix D Clockwise Prompting

### v3

@@ 6.1 Model Bias Towards 0° and 90°
Figure 3: Confusion matrix of true vs. predicted rotations for GPT-4o using CoT prompting, summed across three runs on RotBench-large. Rows represent ground-truth labels, columns represent predicted labels. The matrix highlights a significant confusion specifically between 90° and 270° rotations.
To further elucidate the specific types of rotational errors made by models, we analyze the confusion matrix for GPT-4o. Figure˜3 shows the confusion matrix obtained from summing predictions across three runs. We see GPT-4o predominantly struggles with differentiating between 90° and 270° rotations. Specifically, the model misclassifies 459 instances of 90° images as 270° and 424 instances of 270° images as 90°. Yet, 0° and 180° show significantly fewer misclassifications. This analysis underscores a critical shortcoming in model performance, suggesting either that the vision encoder is not providing sufficient signals to distinguish between clockwise and counter-clockwise rotations, or that the MLLM is not adequately incorporating the visual information into its reasoning.
@@ 6.2 Distinguishing Clockwise from Counter-clockwise Rotations
Table 3: Accuracy of different models in identifying 90° clockwise (CW) versus counter-clockwise rotation (CCW). Column indicates ground truth rotation (GT). A small number of responses are incorrectly identified as being right-side up. These responses are excluded from the table’s statistics.
@@
Following the analysis in Section˜6.1, we examine whether MLLMs can distinguish between clockwise (CW) and counter-clockwise rotations (CCW) by simplifying the four-way classification task into a binary classification task. Using the 90° and 270° images from RotBench-large, we prompt MLLMs to determine whether an image has been rotated 90° CCW or CW. [^6] Note that 90° CW is equivalent to a 270° CCW. By asking models to explicitly differentiate the two directions, we hope to provide a clearer signal for directional understanding. Figure˜4 shows an example question where the GPT-4o provides an incorrect answer.
Results.Table˜3 tests CW vs. CCW classification on GPT-4o and Qwen-2.5-VL-7B-Instruct (we additionally provide similar smaller-scale tests on RotBench-small for GPT-5 in Appendix˜G). Table˜3 indicates that GPT-4o has a significant bias toward identifying rotations as 90° counter-clockwise. GPT-4o correctly identified only 52 out of 300 clockwise rotations, whereas Qwen-2.5-VL-7B-Instruct correctly identified only 23 out of 300 clockwise rotations. The models consistently default to labeling ambiguous or uncertain rotations as counter-clockwise, hinting towards a potential underlying perceptual bias (Appendix˜H). While we find that the most recent model tested, GPT-5, has less bias and is better able to identify clockwise rotations, this improvement is not reflected in Table˜2, where the model performs on par with others; in other words, poor performance on RotBench cannot be solely explained by a model’s ability to distinguish clockwise from counterclockwise rotation. See Appendix˜G for further discussion. These findings strongly indicate that the MLLMs tested have limitations in reliably distinguishing between CW and CCW rotational directions, offering a potential explanation for why distinguishing 90° and 270° CCW rotations presents such a challenging task. Further corroborating this idea, we show that evaluating Qwen-2.5-VL-7B-Instruct, GPT-4o, o3 on RotBench-small using clockwise angles does not result in significant performance differences (Appendix˜F).
@@ 6.3 Can Fine-tuning Solve RotBench?
Setup.  To assess whether specialized training can mitigate these performance issues, we conduct fine-tuning experiments using Qwen-2.5-VL-7B-Instruct. During data filtering, we find some images in Spatial-MM closely resemble each other (Fig.˜11). If we fine-tune on images in Spatial-MM not selected for RotBench, the existence of similar images in the training and testing sets may lead to inflated performance. To prevent such instances of overfitting, we train on 1000 images from MS COCO Lin et al. (2015) and evaluate performance on RotBench-large. Appendix˜D provides further training details.
Results.Fig.˜5 reveals a high and consistent accuracy for 0° throughout training, indicating robust recognition of upright images. Performance on 180° gradually improves, stabilizing around 0.8 after approximately 7000 images. However, accuracies for 90° and 270° exhibit substantial oscillations, suggesting that the model did not achieve stable improvement. The model appears to be caught in a cycle, alternating between accuracy gains and losses for these two rotations. This phenomenon is also reflected in our main results (Section˜5), where using CoT prompting improved accuracy on 270° images at the expense of 90°. The unstable performance may result from potential representational constraints in current visual encoders that limit visual understanding capabilities, particularly regarding subtle rotational distinctions.
6.4 Normalized Rotation Voting
Accuracy (↑)
Model 90° 180° 270°
Qwen-2.5-VL-7B-Instruct
Zero-shot 0.95 0.57 $0.03$ $0.15$
Chain-of-Thought $0.86$ $0.20$ $0.13$ $0.30$
Normalized Rotation Voting $0.54$ $0.52$ 0.56 0.52
GPT-4o
Zero-shot $0.87$ $0.65$ $0.85$ $0.21$
Chain-of-Thought 0.91 $0.59$ 0.86 $0.21$
Normalized Rotation Voting $0.86$ 0.88 $0.80$ 0.86
Algorithm 1 Normalized Rotation Voting
[…]
3:for $i\leftarrow 0$ to |images|1|images|-1 do
4:  $rot\leftarrow\text{call\_model}(images[i])$
5:  $rot\_norm\leftarrow(rot-i\times 90)\bmod 360$
6:  append rot_normrot\_norm to rotationsrotations
7:end for
[…]
Intuition and Setup.  Our results show frontier MLLMs are able to reliably identify 0° and 180° images. Leveraging this pattern, we propose a voting approach to identify image rotation that explicitly exploits this asymmetry in model behavior. Given an input image of unknown orientation, we can further rotate this image 0°, 90°, 180°, and 270°. We separately prompt the model to identify the rotation of each image, then normalize the predictions by subtracting the applied rotation, effectively shifting the angle into a common reference frame (Algorithm˜1). Regardless of the original rotation, two of the further rotated images would correspond to 0° and 180° in the ground-truth reference frame. If the model can correctly identify these two images, there is a high probability that the majority vote would reveal the ground-truth rotation. We evaluate GPT-4o and Qwen-2.5-VL-7B-Instruct on RotBench-small using this normalized majority voting approach to test whether our intuition hold empirically.
Results.Table˜4 shows the accuracy on each image orientation using normalized rotation voting. In general, we see a much more even performance distribution across all four rotations compared to using zero-shot or chain-of-thought prompting. Qwen-2.5-VL-7B-Instruct achieves substantially stronger zero-shot performance on 0° and 90° (0.95 and 0.57) compared to 180° and 270° (0.03 and 0.15). Normalized rotation voting achieves $0.5$ performance on all orientations. GPT-4o achieves $0.8$ or above accuracy on all orientations. In particular, performance on 270° improved substantially compared to zero-shot prompting.
@@ 7 Conclusion
Limitations
Our work shows MLLMs struggle to accurately identify the rotation of input images rotated (0°, 90°, 180°, and 270°). In real-world scenarios, images are typically rotated by arbitrary angles. Therefore, our work may in fact overestimate model capabilities. Moreover, we did not evaluate larger open-weight models due to resource and hardware limitations. Rather, we evaluated the most recent and largest proprietary MLLMs through their APIs. Since these proprietary models dominate open-weight models on nearly all benchmarks (OpenAI, 2025c; Gemini Team, 2025), we expect larger open-weight models to perform worse or on par with the top proprietary models tested.
@@ A.1 Further details on Stage 1 and 2 Filtering
Stage 1.  We randomly sample 300 images from Spatial-MM Shiri et al. (2024). Spatial-MM is divided into two splits, one-subject and two-subject, respectively comprised of images that contain one and two primary subjects. As models may use image dimensions to infer orientation, we crop each image into a square before further processing. We sample 150 images from each split and show the resulting 300-image dataset to the Stage 1 annotator. For each image, the annotator can decide to accept, discard, or flag. The annotator accepts an image if it (1) contains easily identifiable rotational visual cues (e.g., a person standing), and (2) the image has meaningful differences when rotated. Images that do not satisfy these criteria are discarded. Occasionally, some images have subtle visual cues, or require more detailed semantic understanding to correctly perceive orientation. These images tend to have primary subjects that do not vary significantly with rotation, or require integrating background signals to identify rotation. The sample image shown in Fig.˜6 is an example of a flagged image. Such images are flagged by the Stage 1 annotator to proceed to Stage 2. We provide further examples and details of accepted, flagged, and discarded images in Fig.˜7. Note that the Stage 1 annotator is also tasked with ensuring no leaked personal information or offensive content is incorporated into RotBench.
@@
Of the first 300 sampled images, only 27 were flagged. Among the flagged images, only two were eventually rejected. We sample two additional images to replace the discarded images. These images form the 300-image RotBench-large.
@@ A.4 Annotator Interface
Figure˜8 shows the interface provided to Stage 2 annotators. The mapping between letter choice and rotation degree is shuffled from question to question, ensuring the model does not suffer from choice-order bias.
@@ A.5 Sample Images
Fig.˜9 displays the various forms of additional information provided to the model. Fig.˜7 provides examples of an image that is accepted, flagged, and discarded during Stage 1 of data filtering. The accepted image has a clear subject and is easily distinguishable when rotated 0°, 90°, 180°, and 270°. The flagged image is more difficult to identify; however, the slight tilt—as opposed to a directly top-down perspective—still enables accurate judgment of rotation. On the other hand, the discarded image has no meaningful signals to distinguish between the four orientations.
@@ A.6 Further details on flagged images
Flagged images typically involve cases where there is a lack of primary subjects or the primary subject does not offer any distinguishable signals under rotation. These images tend to require deeper comprehension to achieve accurate classification. For instance, Fig.˜6 displays a blank projector canvas above some colorful couches on a patio. The orientation of a projector screen cannot be reliably identified by where its supporting structures are located—the screen can be hung (cable hangs down), pitched up (supporting stand on the ground), or attached to a wall (support frame located sideways of the screen). Therefore, correctly identifying the orientation of the image requires understanding the correct orientation of the couches.
A.7 Motivation for Choosing Spatial-MM
There are multiple ways a model may reason about rotation. If an image has only one primary subject, rotation must be identified using the subject itself and background information. However, if there are multiple primary subjects, the spatial relationship between the two can also provide valuable information. We wish to curate a dataset that balances these two forms of reasoning. As previously mentioned, Spatial-MM’s images already distinguish between these two categories, greatly facilitating this balancing when creating RotBench. Furthermore, Spatial-MM is composed of images scraped from the internet, covering a wide range of lifestyle, portrait, and landscape images. This diversity leads to better generalization to other real-world scenarios.
Appendix B Other Relevant Work
Camera orientation estimation.  Camera orientation estimation is also a well-studied task in computer vision Xian et al. (2019). Instead of predicting the rotation of the image, camera orientation estimation seeks to predict the spatial location of the camera when capturing an image. Contemporary approaches of this task use deep networks to directly predict orientation parameters from image features in an end-to-end manner Xian et al. (2019); Lee et al. (2021, 2020).
Sensitivity to visual perturbations.  Previous work has shown that visual encoders and MLLMs are sensitive to simple image transformations. Anis et al. (2025) evaluate CLIP Radford et al. (2021) and SigLIP Zhai et al. (2023) on a suite of common image transformations—rotations, flips, noise, etc.—revealing substantial gaps between human and model understanding. Usama et al. (2025) finds MLLMs exhibit distinct failure patterns in scene‐text and object reasoning tasks when applying ImageNet-C corruptions Hendrycks and Dietterich (2019) to image inputs. Together, these studies highlight that despite strong clean‐image performance, visual encoders and MLLMs are highly sensitive to photometric and geometric distortions.
Appendix C Generalization and Potential Downstream Applications
While our work provides a detailed analysis into the task of identifying image rotation, showing that such a simple and intuitive task for humans remains challenging for frontier reasoning MLLMs. Despite its straightforward nature, image rotation estimation requires strong perception, spatial reasoning abilities, and general world knowledge. For each image, the model must recognize which objects are relevant to rotational understanding, identify their spatial relationships, and determine whether there exists any irregularities in such relationships. Furthermore, computer-vision tasks typically tend to emphasize the primary subject, or subjects at the forefront of an image (e.g., object extraction, counting, segmentation, etc). However, image rotation often requires the model to also interpret subtle background information.
These abilities are a fundamental prerequisite for higher-level spatial reasoning. Our findings expose systematic blind spots which can lead to failures in more complex downstream tasks, such as understanding relative positions (“left/right/above/below”) or reasoning under rotated cameras.
In addition, we further elaborate on the two downstream applications outlined in Section˜1:
  • Robotics: AI tasks in manufacturing may require cameras mounted on robotic arms that can move in multiple axes. Researchers interested in training a robotic arm to perform complex maneuvers – such as orienting a drill bit to reach awkward angles – with a high degree of autonomy, would require a model that can accurately identify the current rotation of the arm.
  • First-person Point-of-View (PoV) footage analysis: Analyzing footage from various extreme sports – such as cliff diving or sky diving – and aerial vehicles – such as camera drones – involves complex rotational movement across multiple axes of rotation. Researchers interested in training RL agents in such environments require the agent to exhibit strong spatial reasoning abilities, among which is the prerequisite ability to identify its current rotation from the image.
Appendix D Model Inference and Training Details
@@
Appendix E Additional Results
We provide experimental results from evaluating Llama-3.2-11B-Instruct on RotBench-large (Table˜5), as well as o3, Gemini-2.5-Pro (Table˜6),Qwen-2.5-VL-3B-Instruct, Claude-4.1-Opus, Phi-4-MM-Instruct, Gemma-3-12B-Instruct, and Molmo-7B-O on RotBench-small (Table˜7. Llama-3.2-11B-Instruct seems to be the weakest model tested, showing inferior performance across all rotations. Without CoT, the model only achieves roughly 0.3 accuracy. The two frontier reasoning models (o3 and Gemini-2.5-Pro) show no significant performance improvement when prompted with auxiliary information. However, they respond positively to rotation grids. Gemini-2.5-Pro achieves the highest accuracy on 90° and 270° (both 0.7) using guided rotation grids, while o3 achieves its highest accuracy on 180° (0.83) using the rotation grid. These results indicate that strong reasoning models are able to utilize the grid to aid them in making a prediction, while weaker models struggle with it.
@@ Appendix E Additional Results
Accuracy on Different Degrees of Rotation
Model 90° 180° 270°
Qwen-2.5-VL-3B-Instruct
Zero-shot $\textbf{0.60}{\scriptscriptstyle\pm 0.00}$ $\textbf{0.52}{\scriptscriptstyle\pm 0.00}$ $0.16{\scriptscriptstyle\pm 0.00}$ $0.22{\scriptscriptstyle\pm 0.00}$
     + Chain-of-Thought $0.18{\scriptscriptstyle\pm 0.00}$ $0.44{\scriptscriptstyle\pm 0.00}$ $\textbf{0.42}{\scriptscriptstyle\pm 0.00}$ $\textbf{0.28}{\scriptscriptstyle\pm 0.00}$
Phi-4-$M$-Instruct
Zero-shot $\textbf{0.86}{\scriptscriptstyle\pm 0.00}$ $0.10{\scriptscriptstyle\pm 0.00}$ $26{\scriptscriptstyle\pm 0.00} 0.0$$02{\scriptscriptstyle\pm 0.00}$
     + Chain-of-Thought $0.34{\scriptscriptstyle\pm 0.00}$ $\textbf{0.20}{\scriptscriptstyle\pm 0.00}$ $\textbf{0.40}{\scriptscriptstyle\pm 0.00}$ $\textbf{0.28}{\scriptscriptstyle\pm 0.00}$
Gemma-3-12B-Instruct
Zero-shot $\textbf{0.96}{\scriptscriptstyle\pm 0.02}$ $0.18{\scriptscriptstyle\pm 0.03}$ $\textbf{0.56}{\scriptscriptstyle\pm 0.02}$ $0.18{\scriptscriptstyle\pm 0.01}$
     + Chain-of-Thought $0.34{\scriptscriptstyle\pm 0.00}$ $0.20{\scriptscriptstyle\pm 0.04}$ $0.40{\scriptscriptstyle\pm 0.01}$ $\textbf{0.28}{\scriptscriptstyle\pm 0.05}$
Molmo-7B-O
Zero-shot $\textbf{1.00}{\scriptscriptstyle\pm 0.00}$ ${\scriptscriptstyle\pm 0.00} 0.00±0.00$$\scriptscriptstyle\pm 0.00$} $0.00{\scriptscriptstyle\pm 0.00}$
     + Chain-of-Thought $0.58{\scriptscriptstyle\pm 0.00}$ $\textbf{0.28}{\scriptscriptstyle\pm 0.00}$ $\textbf{0.52}{\scriptscriptstyle\pm 0.00}$ $\textbf{0.32}{\scriptscriptstyle\pm 0.00}$
Claude-4.1-Opus
Zero-shot $0.96{\scriptscriptstyle\pm 0.00}$ $\textbf{0.38}{\scriptscriptstyle\pm 0.00}$ $\textbf{0.48}{\scriptscriptstyle\pm 0.00}$ $\textbf{0.30}{\scriptscriptstyle\pm 0.00}$
     + Chain-of-Thought $\textbf{0.98}{\scriptscriptstyle\pm 0.00}$ $0.28{\scriptscriptstyle\pm 0.00}$ $0.42{\scriptscriptstyle\pm 0.00}$ $0.12{\scriptscriptstyle\pm 0.00}$
Table 7: Average classification accuracy under different image rotation angles (0°, 90°, 180°, 270°) for Qwen-2.5-VL-3B-Instruct, Phi-4-MM-Instruct, Gemma-3-12B-Instruct, Molmo-7B-O, Claude-4.1-Opus and various forms of auxiliary information across a single run on RotBench-small. Accuracy is scored on a four-way classification task.
Appendix F Clockwise Prompting